# EmoFace: Emotion-Content Disentangled Speech-Driven 3D Talking Face with Mesh Attention

### v1

@@ Methods
In addition, we compare the difference in effectiveness between the Mesh Attention module and the Add operation. Table 2 also demonstrates that directly adding up the gains predicted by the two branches without Mesh Attention leads to a great degradation of model performance, whereas a weighted sum seems more reasonable. We also provide a comparison demo in the Supplementary Video. We believe that although the lips are directly controlled by the content of the speech, the gain from emotion is also quite important. Hence, we visualize the attention weight of the emotion branch with darker colour representing larger weights, as shown in Figure 5. It is obvious that emotion brings great effects on the amplitude and direction of movements in the mouth, jaw and eye regions. In other words, these regions are strongly correlative to emotion.
In the Supplementary Video, we visualize the animation demo of the content branch. It performs normally in non-silent frames but shows obvious lip jitter in silent frames, thus also resulting in unstable output of the Addition. However, EmoFace dynamically combines the two branches with Mesh Attention to obtain better animation results in terms of lip synchronization, expression synchronization, movement amplitude and coherence.
@@
User study
@@
We design a comprehensive research questionnaire to evaluate the effectiveness of EmoFace and compare it with FaceFormer, CodeTalker, EmoTalk and SelfTalk. Apart from the other methods, EmoFace can extract emotional information from speech. So we provide 12 sets of comparison results with 8 emotions, and finally make 48 questions targeting four aspects: full-face, lip synchronization, eye movement and emotion expression. The questionnaire shows comparison videos to the respondents and asks them to rate the effectiveness of each algorithm according to its realism. We count the Mean Opinion Score (MOS) of all methods and EmoFace has a higher MOS compared to other approaches. More specifically, as shown in Table 3, our full-face animation receives the most positive feedback. Meanwhile, we find that SelfTalk and EmoFace score very close to each other in terms of lip synchronisation, while FaceFormer and EmoTalk are slightly worse. EmoFace shows a significant advantage over other methods in emotional expression and in eye movements. In conclusion, our method outperforms other methods in the questionnaire study.
@@
Conclusions and Discussion
In this paper, we propose EmoFace, an emotional speech-driven 3D talking face model. Our model disentangles the emotion and content of the speech through two independent branches, predicting the mesh offsets driven by emotion and content, respectively. In order to further improve the accuracy of the prediction, we design new Mesh Attention to integrate the output mesh offset of the two branches. In particular, a novel graph-based SpiralConv3D is adopted to fuse spatiotemporal features of mesh sequences. To address the problem of lacking 3D emotional talking face data, we construct a high-quality 3D emotional talking face dataset 3D-RAVDESS that contains mesh vertices. We train EmoFace under our proposed self-growing scheme and conduct extensive experiments to demonstrate that our model outperforms existing state-of-the-art methods and receives better user experience feedback. Meanwhile, EmoFace can adapt a non-silent frame as the initial input frame, which provides great potential for future adaptation to more diverse datasets.
Although our model gets state-of-the-art results, there are still some limitations to address in future work. On the one hand, speech driven methods do not model expressions and motions that are unrelative to audio, for example, eye blinks. Our next work will take the video input into account to obtain better animation results. On the other hand, the Mesh Attention proposed in our work meets large amount of computation because of the 3D spiral convolution operator. More explorations will be done to lower the computation costs.

### v3

@@ 4.4 Ablation experiments
4.4.1 Impact of self-growing scheme.
We train our method using teacher-forcing without self-growing and obtain higher LVE and EVE, as shown in Table 3. We believe the reason is that guiding is too strong, leading to poor robustness, and the prediction errors in the previous frames accumulate and affect the subsequent frames. However, with a gradually weakened guiding in the self-growing scheme, EmoFace is trained to take the errors of the previous frames into account, which is similar to the situation during inference.
4.4.2 Impact of Mesh Attention.
We compare the difference in effectiveness between the Mesh Attention module and the Add operation. Table 3 also demonstrates that directly adding up the gains predicted by the two branches without Mesh Attention leads to a great degradation of model performance, whereas a weighted sum seems more reasonable. This indicates that the contribution of the two branches differs in the different head regions.
4.4.3 Hyperparameters of SpiralConv3D.
A series of ablation experiments on dilation, kernel size and time duration are provided, as shown in Table 4. In terms of dilation coefficients, the third to fifth rows show that proper dilation can improve performance, while excessive dilation leads to performance degradation. SpiralConv3D may under-consider nearby vertex features with too large dilation coefficient. Furthermore, with respect to kernel size, the first to third rows indicate that large convolution kernels bring high generation quality. SpiralConv3D with a larger kernel size integrates vertex features within a larger spatial receptive field, which has a significant effect on the performance improvement. Finally, the fourth and the last two rows demonstrate the importance of large temporal receptive field. A larger time duration means that the connectivity relationship of the vertices between more video frames are considered. Time duration coefficient of 1 degrades SpiralConv3D to SpiralConv2D, which only spatially fuses the features and performs worse than configuration of the fourth row. Based on the results of the quantitative comparison, we adopt the configuration of the fourth row.
Dilation
Kernel
Size
Time
Duration
LVE
(×$\times 10^{-5}$mm)
EVE
(×$\times 10^{-5}$mm)
1 5 3 5.3472 1.1495
1 9 3 5.2177 1.1089
1 18 3 4.7527 1.0244
2 18 3 4.8863 0.9509
4 18 3 5.1323 1.1390
2 18 1 5.3476 1.1715
2 18 2 5.5372 1.2144
Table 4: Ablation study of SpiralConv3D.
@@
4.5 Visualization
We believe that although the lips are directly controlled by the content of the speech, the gain from emotion is also quite important. Hence, we visualize the attention weight of the emotion branch with darker color representing larger weights, as shown in Figure 4. It is obvious that emotion brings great effects on the motion amplitude and direction in the mouth, jaw and eye regions. In other words, these regions are strongly correlative to emotion.
The Supplementary Video shows the animation of the content branch. It performs normally in non-silent frames but shows obvious lip jitter in silent frames, which also leads to unstable Addition output. However, EmoFace dynamically combines two branches with Mesh Attention to obtain better animation results in terms of lip synchronization, expression synchronization, movement amplitude and coherence. More results under VOCASET-Test, 3D-RAVDESS-Test and long speech in the wild, performance on different emotions and comparative demonstrations of ablation experiments are shown in the Supplementary Video.
5 Conclusion
In this paper, we present EmoFace, an emotional speech-driven 3D talking face model. Our model disentangles the emotion and content from the speech and predicts the mesh offsets driven by emotion and content, respectively. To further improve the prediction accuracy, new Mesh Attention integrates the output mesh offsets. Particularly, a novel graph-based SpiralConv3D is adopted to fuse spatio-temporal features of mesh sequences. Extensive experiments conducted under self-growing scheme using both public dataset VOCASET and newly constructed high-quality dataset 3D-RAVDESS demonstrate that our model outperforms existing state-of-the-art methods and receives better user experience feedback.
Limitations
Although our model gets state-of-the-art results, there are still some limitations to address in future work. On the one hand, speech driven methods do not model expressions and motions that are unrelative to audio, for example, eye blinks. Our next work will take the video input into account to obtain better animation results. On the other hand, the Mesh Attention proposed in our work meets large amount of computation because of the 3D spiral convolution operator. More explorations will be done to reduce the computation costs.
Ethical Statement
The existing scientific artifacts used in this work are consistent with their intended use and our work is for research purposes only and should not be used outside of research contexts. In addition, since face data can be used for generating content that may jeopardize privacy, we must act responsibly by considering the aspects related to privacy and ethics.
Appendix A Appendix
A.1 Training Details
Our method takes speech data as input and also provides mesh templates, emotion levels, and speaker IDs as conditions. The sampling rate of the speech is $16kHz$ and the frame rate of the mesh sequence is 30 frames per second. During training, the model is end-to-end optimised using the Adam optimizer. The learning rate and batch size are set to $10^{-4}$ and $1$ respectively. The model is trained on a single NVIDIA A100 with 200 epochs on VOCASET and 3D-RAVDESS.
A.2 Emotion-Content Disentanglement Module
We set a pseudo-training pair of samples with the same content but different emotions. The emotion features and content features are extracted, respectively, and the two emotion features are exchanged as input of transformer decoder to implement cross reconstruction, as shown in Figure 5. The entire architecture of EmoFace is described in detail as follows:
$F_{c1}=E_{c}(A_{c1,e1}),$ (12)
$F_{e1}=E_{e}(A_{c1,e1}),$ (13)
$F_{e2}=E_{e}(A_{c1,e2}),$ (14)
$M^{c}_{c1,e1}=M^{c}_{c1,e2}=D_{c}(F_{c1}),$ (15)
$M^{e}_{c1,e1}=D_{e}(F_{e1}),M^{e}_{c1,e2}=D_{e}(F_{e2}),$ (16)
$M_{c1,e1}=\epsilon(M^{e}_{c1,e1},M^{c}_{c1,e1}),$ (17)
$M_{c1,e2}=\epsilon(M^{e}_{c1,e2},M^{c}_{c1,e2}),$ (18)
where $\epsilon$ is the Mesh Attention module, $A_{cx,ey}$ represents audio with content $x$ and emotion $y$, $M_{cx,ey}$ represents the mesh sequence with content $x$ and emotion $y$, $F_{c}$ and $F_{e}$ denote the content features and the emotional features, respectively, $E$ and $D$ are the encoder and decoder respectively.
Figure 5: Supervision training strategy of emotion-content disentanglement module. Various inputs of speech, conveying same content and different emotions, are processed to generate cross-reconstructed mesh vertex offsets representing distinct combinations of facial expressions. Supervisions are added to both two branches and the final output.
A.3 Baseline Methods
Figure 6: FLAME Head Templates for 24 actors in 3D-RAVDESS.
Our method is compared with FaceFormer Fan et al. (2022), CodeTalker Xing et al. (2023), EmoTalk Peng et al. (2023b), SelfTalk Peng et al. (2023a) and TalkingStyle Song et al. (2024) and achieves state-of-the-art performance on both VOCASET Cudeiro et al. (2019) and 3D-RAVDESS datasets. Since the official EmoTalk is a blendshape-based method that is incompatible with the marked facial mesh vertices provided by the datasets, we modify its output layer to directly predict the offsets of the vertices. Our method, FaceFormer, CodeTalker and TalkingStyle require conditions on a training speaker identity during inference. Therefore, for unseen subjects in the test dataset, we follow FaceFormer and CodeTalker to obtain the predictions by conditioning on training identities. To adapt the input of non-silent initial frame to the 3D-RAVDESS dataset, we use the first frame as a condition in the training of EmoFace, FaceFormer, CodeTalker and TalkingStyle. All of these models are autoregressive, facilitating the use of input data from the first frame, while SelfTalk and EmoTalk are not autoregressive models. For VOCASET, we use the pre-trained models provided by the five methods to evaluate the vertex error on the VOCA-Test. For 3D-RAVDESS, we retrain other methods under the official experimental configuration. During the evaluation, we compute the error directly between the output vertices and the ground truth. It is worth noting that since CodeTalker training on 3D-RAVDESS crashes no matter how the hyperparameters are adjusted, as mentioned in DF-3DFace Park et al. (2023) and LaughTalk Sung-Bin et al. (2024). Thus, we follow these two previous work to directly use the pre-trained model trained on VOCASET for zero-shot prediction.
Besides, The 3D-ETF dataset used in EmoTalk is a blendshape dataset and official EmoTalk has to convert the predict blendshape coefficients into mesh vertices. However, EmoTalk has not yet released its converter. Thus, we modified EmoTalk’s output layer to directly predict vertex offsets to adapt to our vertex dataset 3D-RAVDESS to make a fair comparison.
A.4 Difference between self-growing and scheduled sampling
The primary distinction lies in how Stage 2 inputs are constructed. Scheduled sampling employs a static mixing strategy (Softmax, Gumbel-Softmax, or Sparemax) to blend ground-truth and Stage 1 predicted frames throughout training. In contrast, our self-growing scheme dynamically transitions through three optimized phases (Eq.6), each designed to address specific training challenges.
Phase 1: Warm-up Period. The model trains exclusively on groundtruth inputs, avoiding disruptions by unreliable early predictions. This period ensures stable initial learning.
Phase 2: Adaptive Prediction Integration. A cosine annealing scheduler gradually increases the proportion of Stage 1 predictions used as inputs. As prediction accuracy improves with training, the adoption rate accelerates (reflected by the increasing derivative of the cosine function in Eq.6), enabling smooth transition to prediction-dominated inputs.
Phase 3: Full Prediction Mode. The system operates solely on Stage 1 predictions, rigorously stress-testing the model against error accumulation – a critical scenario omitted by scheduled sampling.
This phased approach ensures robust training, while scheduled sampling lacks both initial stabilization and final robustness validation phases.
A.5 Experiments between self-growing and autoregression
We have compared our method with autoregression in our experiments. After extensive hyperparameter tuning failed to achieve convergence in autoregressive training, we excluded these results from Table 3. The non-convergence may be caused by two key factors: (1) the inherent complexity of 3D-RAVDESS dataset samples, which exhibit high variability in emotional expression dynamics, and (2) the error accumulation characteristic of autoregression, where initial prediction inaccuracies amplify catastrophically during sequential generation.
A.6 3D-RAVDESS Dataset
In this study, we reconstruct 3D face mesh sequences from 2D videos based on FLAME topology to construct a large 3D emotional talking face dataset, 3D-RAVDESS.
Specifically, 1440 videos from the RAVDESS Livingstone and Russo (2018) are processed by converting them into 30 frames per second and capturing the mesh vertices for each frame by EMOCA Daněček et al. (2022). To enhance the quality of the dataset and reduce frame-to-frame jitter, a Kalman Filter is applied to the output FLAME parameters, which significantly improves the smoothness of facial animation. The 3D-RAVDESS dataset generates 159,702 frames of mesh vertices, which amounts to approximately 1.5 hours of video content.
Furthermore, 3D-RAVDESS dataset contains 24 templates for different actors (as shown in Figure. 6). The actors all read sentences to express specific emotions, including neutral, calm, happy, sad, angry, fearful, disgusted, and surprised. A supplementary video demonstrates the animation results of our EmoFace under speech with different emotional types and intensities.
3D-RAVDESS outperforms the dataset used in EMOTE with its strictly controlled multimodal synchronization and standardized actor performances, making it more reliable for precise affective computing research.
Figure 7: Rendering Process.
A.7 User study
@@
We design a comprehensive research questionnaire to evaluate the effectiveness of EmoFace and compare it with FaceFormer, CodeTalker, EmoTalk and SelfTalk. We provide 12 sets of comparison results on 3D-RAVDESS-Test and VOCA-Test with 8 emotions, and finally make 48 questions targeting four aspects: full face realism, lip synchronization, eye movement and emotion expression. The questionnaire shows comparison videos to the respondents and asks them to rate the effectiveness of each algorithm. We calculate the Mean Opinion Score (MOS) of all methods and EmoFace obtains the highest MOS, as shown in 5, suggesting that our method receives the most positive feedback.
@@ Methods
A.8 Application
The final render results with textures can be obtained based on the output mesh of our method. The rendering process of transforming a 3D mesh into a final rendered image is shown in Figure. 7. The requirements include materials, textures and scene setup, each contributing to the visual fidelity and realism of the final output.
A.8.1 Materials and Textures
Materials and textures are applied to the mesh to define its visual properties. On the one hand, materials define surface characteristics such as color, reflectivity and transparency, which determine how the mesh interacts with light. On the other hand, textures add details like patterns and surface irregularities. In our work, we use AlbedoGAN Rai et al. (2024) to obtain an albedo map image, a texture image that represents the color information without any lighting effects.
A.8.2 Scene Setup
The scene setup involves placing the mesh within a virtual environment along with cameras, lighting and shading. Camera setup defines the camera’s position, orientation, and properties to capture the scene appropriately. And lighting adds light sources to illuminate the scene, which significantly impacts the final appearance of the mesh. The shading calculates the color and intensity of each pixel based on lighting and material properties.
A.8.3 Rendering
A differentiable renderer is applied for high-quality renderings from the mesh and the albedo map.
A.9 Supplementary Video
We provide a Supplementary Video showing the results under VOCASET-Test, 3D-RAVDESS-Test and long speech in the wild, performance on different emotions and comparative demonstrations of ablation experiments.